OpenStudy (anonymous):

Write the equation of a circle with a center at (5, 5) and a diameter of 13.

OpenStudy (anonymous):

(x – 5)^2 + (y – 5)^2 = 42.25 (x + 5)^2 + (y + 5)^2 = 169 (x + 5)^2 + (y + 5)^2 = 13 (x + 5)^2 + (y + 5)^2 = 42.25

OpenStudy (anonymous):

it has to be between a and c

OpenStudy (barrycarter):

Hint: what's the distance formula (Pythagorean Theorem)?

OpenStudy (anonymous):

not sure.. :/

OpenStudy (cwrw238):

number on the right = r^2 (radius squared) and radius = half diameter

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

soo 13^2?

OpenStudy (cwrw238):

noo - 13 = diameter

OpenStudy (anonymous):

so half of 13

OpenStudy (anonymous):

6.5?

OpenStudy (cwrw238):

general equation of circle is (x-a)^2 + (y-b)^2 = r^2 yes r = 6.5

OpenStudy (anonymous):

thats not one of the choices though

OpenStudy (cwrw238):

(a,b) is center

OpenStudy (cwrw238):

r^2 not r

OpenStudy (cwrw238):

6.5^2

OpenStudy (anonymous):

ok! i see....so it would be d not a correct?

OpenStudy (anonymous):

@cwrw238 ?

OpenStudy (cwrw238):

its (x-a)^2 right? and a = 5 so thats (x-5) is that a or d?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks @cwrw238

OpenStudy (cwrw238):

yes